Given that f(x) = (4 − x2)^1/2 and h(x) = (3 − x)^1/2, Find foh(x).
1
Expert's answer
2020-07-09T19:44:24-0400

if x2 means x^2 then

"foh(x) = f(h(x))= \sqrt{4 - (\sqrt{3-x})^2} = \sqrt{4-(3-x)}=\sqrt{1+x}"

if x2 means x*2 then

"foh(x)=\sqrt{4-(\sqrt{3-x})*2}"